# Analytical and Numerical Studies of Dark Current in Radiofrequency Structures for Short-Pulse High-Gradient Acceleration

### v2

@@ II.3 SEY under multipacting resonance
We now turn to the evaluation of SEY under multipacting resonance conditions. Multipacting resonance occurs when secondary electrons emitted from a surface are accelerated by an RF field such that their time of flight $\tau$ is synchronized with the RF period $T$ [46, 38]. The resonance order $N$ is given by $N=\tau/T$. When the SEY associated with a given resonance exceeds unity, the number of secondary electrons can increase with each RF cycle, potentially leading to an electron avalanche during a long RF pulse.
To identify resonance conditions, we begin with a specified initial RF phase $\phi_{0}$ and axial emission position $z_{0}$, and calculate the electron trajectory over multiple iterations. A convergence criterion is applied, requiring both the RF phase $\phi$ at which the electron leaves the surface and the axial position $z$ to stabilize within a tolerance of $1\times 10^{-6}$ over ten consecutive iterations.
Figure 7: Multipacting susceptibility plots at resonance showing the SEY $\delta$ at various $E_{z0}$ with respect to (a) resonant axial emission position $z_{\text{res}}$, (b) resonant emission angle $\alpha_{\text{res}}$, and (c) resonant emission RF phase $\phi_{\text{res}}$.
Figure 8: Simulated electron trajectories emitted from (a) the iris and (b) the photocathode surface, under a peak axial electric field amplitude of 325 MV/m at the center of the photocathode surface. Color-coded lines represent trajectories of electrons launched from different initial positions.
Figure 6 shows trajectories of multipacting resonance modes in the $\Delta r\text{--}z$ plane for various resonance orders $N$. SEY values are calculated for each mode across a range of initial conditions. The corresponding susceptibility map in Fig. 7 plots SEY $\delta$ as a function of $E_{z0}$ and three key parameters: (a) resonant axial position $z_{\text{res}}$, (b) emission angle $\alpha_{\text{res}}$, and (c) emission phase $\phi_{\text{res}}$. These plots reveal distinct bands of sustained multipacting. As $E_{z0}$ increases, dominant resonance orders shift from higher ($N=7$) to lower ($N=1$), a typical trend in RF multipactor behavior. Identifying these bands provides insight into multipacting onset and informs suppression strategies.
@@ III Dark Current Simulations
To investigate dark current behavior under short-pulse RF excitation, we perform numerical simulations using the cst Particle-In-Cell (PIC) and particle tracking solvers. This complements the steady-state analytical model by capturing transient effects in realistic cavity geometries. We begin by modeling field emission from high-gradient regions, such as the cavity iris and photocathode. These field-emitted electrons can act as seed particles for multipacting-driven growth and subsequent dark current buildup, as discussed later.
@@ III.1 Field Emission Simulations
Field emission is a critical process in RF cavities, where strong local electric fields can extract electrons from metal surfaces. These electrons may disrupt the primary beam [50, 51], trigger RF breakdown [29], or initiate dark current through multipacting. Extensive efforts have been made to model and characterize field emission across various RF structures [31, 52, 27, 53].
Field emission simulations were performed using the cst PIC solver to investigate the low dark current observed during RF conditioning of the $X$-band photogun cavities. As discussed in Sec. I, no dark current signal was detected above the noise threshold [5]. This can be attributed to two factors: (1) the short RF pulse limits the total emitted charge, and (2) emitted electrons must traverse a long, narrow beam pipe (the inner conductor of the coaxial coupler) to reach diagnostics [54]. This is illustrated in Fig. 8: in (a), most electrons emitted from the iris strike nearby surfaces; in (b), only electrons emitted near the center of the photocathode and within a narrow angular range are able to exit the beam pipe. Simulations estimate that only about $10^{-5}$ of the total electrons emitted from the iris reach the end of the beam pipe. The electrons that strike cavity surfaces may initiate multipacting when resonance conditions are met.
@@
Figure 10: Multipacting resonance trajectories simulated in the full cell (defined in Fig. 9). The corresponding axial field amplitudes $E_{z0}$ for resonance orders $N=1\text{--}7$, evaluated at point $P_{3}$ (the center of the full cell), are 106, 61, 39, 31, 26, 22, and 19 MV/m, respectively. These trajectories show good agreement with the analytical results in Fig. 6, as the RF field distribution in the full cell is well approximated by the analytical model.
Figure 11: Multipacting resonance trajectories in the half-cell for (a) Region I and (b) Region II. In Region I, the axial fields $E_{z0}$ evaluated at point $P_{1}$ for resonance orders $N=1\text{--}4$ are 112, 61, 39, and 28 MV/m, respectively. In Region II, the axial fields $E_{z0}$ evaluated at point $P_{2}$ for resonance orders $N=1\text{--}5$ are 95, 46, 30, 25, and 21 MV/m, respectively.
III.2 Multipacting simulations in the short-pulse regime
@@
In particle tracking simulations, electron trajectories were analyzed for emission from the sidewalls of the three regions as defined in Fig. 9, the full cell, and Region I and Region II in the half cell. Sample trajectories satisfying multipacting resonance conditions are presented in Figs. 10 and 11. The resonant trajectories and the associated electric field amplitudes in the full cell (Fig. 10) exhibit excellent agreement with the analytical results shown in Fig. 6 due to the similarity in the RF field pattern, providing further validation of the analytical model.
Figure 12: Temporal evolution of the electric field amplitudes and the number of multipacting electrons for (a) Region I, (b) Region II, and (c) the full cell. In each subplot, the on-axis electric field at point $P_{1}$, $P_{2}$, or $P_{3}$ (red solid line), the electric field across the photocathode surface (black dashed line), and the number of multipacting electrons (blue solid line) are shown. For Region I, the dark current peaks corresponding to $N=1\text{--}4$ appear at 113, 66, 43, and 30 MV/m; for Region II, the peaks for $N=1,2,4,5$ are at 100, 47, 28, and 20 MV/m; and for the full cell, the peaks for $N=1,2,3,5$ are at 113, 51, 35, and 27 MV/m. Higher-order resonance modes ($N\geq 6$) also contribute to later peaks in the full cell.
@@
In PIC simulations, we use again the short input RF pulse with a 3 ns rise, 3 ns flat top, and 3 ns fall and a peak power of 200 MW. Seed electrons were introduced following the method in Ref. [38]. Secondary emission was modeled using the Vaughan formalism [49], with a maximum yield $\delta_{max}=2.1$ at an impact energy $W_{i}=150$ eV. Figure 12 shows the temporal evolution of the electric field and the number of multipacting electrons in the three regions. Peaks in the electron population correspond to distinct multipacting resonance orders, and the gradient at resonance are in strong agreement with the multipactor resonance trajectory analysis, confirming that electron growth is driven by multipacting resonances.
Figure 14: Temporal evolution of the electric field (red) and number of multipacting electrons (blue) using a longer RF pulse with a 30 ns rise.
To further confirm that dark current growth arises from multipacting resonances, we varied the peak power of the input RF pulses while maintaining a fixed temporal profile. The resulting time evolution of the number of multipacting electrons is shown in Fig. 13. For a given cavity geometry, the field gradients at which multipacting resonance occurs are fixed. Adjusting the input power changes the time at which these gradients are reached, thereby shifting the onset of the resonance modes. The dark current peaks consistently occur at the same axial field amplitudes across different power levels, matching the resonance orders identified earlier. Each mode persists for a duration that depends on the field ramp rate. At lower gradients, the field ramp rate is reduced, so the peak dark current is higher.
These analyses show that short RF pulses can reduce dark current. Fig. 14 shows the simulation results with longer RF pulses as a comparison, where the gradual rise in field amplitude allows resonance conditions to persist longer, resulting in rapid growth of the multipacting electron population. In contrast, short RF pulses sweep through each resonance condition more quickly, limiting the duration over which the field remains within the amplitude window required for sustained multipacting. As a result, the buildup of secondary electrons is effectively suppressed. This comparison demonstrates that the short-pulse regime reduces dark current by limiting the time available for resonance-driven electron multiplication.
@@ IV Impact of dark electrons
As dark electrons accumulate through field emission and multipacting, electron clouds can form. To observe this process, we performed cst PIC simulations with the cavity operated at the multipacting resonance conditions identified above. Seed electrons were briefly introduced near the cavity sidewall to initiate secondary emission. When the field satisfied resonance conditions, the electron population grew exponentially, forming a localized cloud near the sidewall.
Figure 15: Normalized reflected RF power as a function of time for different plasma electron densities $n_{e}$, modeled as a 0.2 mm thick layer on the cavity sidewall. All cases use the same short RF input pulse. Higher plasma densities lead to further reductions in the reflected power.
A critical consequence of multipacting-induced electron clouds is dark-current-induced beam loading. To model this effect, we approximate the electron cloud as a cold plasma layer with a thickness of 0.2 mm, located along the cavity sidewall. The plasma response is described using Drude dispersion, appropriate for an electron-only, collisionless medium. The effective permittivity of the layer is determined by an assumed plasma electron density $n_{e}$. Figure 15 shows the normalized reflected RF power for different plasma densities. As $n_{e}$ increases, the reflected power decreases due to greater energy absorption by the plasma [55]. This behavior is consistent with the experimentally observed drop in reflected power during conditioning of the X-band photocathode cavities (see Fig. 2).
@@ V Conclusion and discussion
Recent advances in short-pulse acceleration have opened new opportunities for achieving ultra-high gradients, but they also demand a reexamination of the underlying RF breakdown mechanisms on previously underexplored timescales. Traditional breakdown studies, grounded in long-pulse operation and steady-state field assumptions, fall short in capturing the transient physics unique to nanosecond-long RF pulses.
In this work, we use analytical and numerical methods to show that RF pulses on the order of a few nanoseconds can effectively suppress dark current growth by disrupting the timing required for multipacting resonance. The analytical model identifies clear resonance bands, while time-domain simulations reveal how rapid field ramping shortens the window for sustained development of multipacting resonances. These results provide strong evidence that short-pulse operation alters the onset dynamics of dark current and RF breakdown. Our simulation results are also consistent with experimental observations of low dark current and beam loading in the $X$-band photogun cavities at AWA.
To advance the understanding of RF breakdown in the short-pulse regime, future work should focus on modeling the dynamics of plasma formation and its interaction with time-varying RF fields, ideally using first-principles simulation tools. Conventional breakdown models often treat plasma evolution as a multi-stage process, with nanosecond-scale electron dynamics coupled to slower changes in surface morphology. However, in the short-pulse regime, where the electromagnetic fields vary rapidly in time, these transient interactions may dominate the breakdown initiation process. Investigating these fast-timescale effects will be essential for refining our understanding of the coupled physics in RF breakdown, and short-pulse experiments provide a promising platform for isolating and studying these mechanisms.

### v3

@@ II.3 SEY under multipacting resonance
We now turn to the evaluation of SEY under multipacting resonance conditions. Multipacting resonance occurs when secondary electrons emitted from a surface are accelerated by an RF field such that their time of flight $\tau$ is synchronized with the RF period $T$ [46]. The resonance order $N$ is given by $N=\tau/T$. When the SEY associated with a given resonance exceeds unity, the number of secondary electrons can increase with each RF cycle, potentially leading to an electron avalanche during a long RF pulse. Here we focus on converged resonant trajectories, i.e., trajectories that are identical across successive iterations. Identification of these resonant trajectories constitutes one original contribution of the analytical theory reported here, relative to Ref. [38]. These converged resonant trajectories are phase-locked and therefore more likely to contribute to dark current growth.
To identify resonance conditions, we fix one parameter, such as the emission angle $\alpha_{0}$, and calculate the electron trajectory while sweeping the initial RF phase $\phi_{0}$ and axial emission position $z_{0}$ until convergence is reached. Convergence is defined by the return RF phase $\phi$, axial position $z$ and the SEY $\delta$ stabilizing within $1\times 10^{-6}$ over ten consecutive iterations. An example of this convergence search process is presented in Appendix B.
Figure 7: Multipacting susceptibility plots at resonance showing the SEY $\delta$ at various $E_{z0}$ with respect to (a) resonant axial emission position $z_{\text{res}}$, (b) resonant emission angle $\alpha_{\text{res}}$, and (c) resonant emission RF phase $\phi_{\text{res}}$. Each panel is a 2D projection of the same resonance islands, with the unplotted variables varying self-consistently to satisfy the resonance conditions.
Figure 6 shows trajectories of multipacting resonance modes in the $\Delta r\text{--}z$ plane for various resonance orders $N$. SEY values are calculated for each mode across a range of initial conditions. The corresponding susceptibility map in Fig. 7 plots SEY $\delta$ as a function of $E_{z0}$ and three key parameters: (a) resonant axial position $z_{\text{res}}$, (b) emission angle $\alpha_{\text{res}}$, and (c) emission phase $\phi_{\text{res}}$. Figure 7(a)-(c) show different 2D cuts of the same resonance islands in the four-dimensional space of ($z_{\text{res}}$, $\alpha_{\text{res}}$, $\phi_{\text{res}}$, $E_{z0}$). In each panel, the other variables adjust self-consistently to satisfy the resonance conditions, rather than being held fixed. These plots reveal distinct bands of sustained multipacting. As $E_{z0}$ increases, dominant resonance orders shift from higher ($N=7$) to lower ($N=1$), a typical trend in RF multipacting behavior. We note that, although these resonance bands shift slightly with the assumed initial energy of the secondary electrons (2 eV in the analytical model here), the range of field gradients that support multipacting resonances remains similar. This is also confirmed by the gradients at which multipacting current growth is predicted in the PIC simulations described in Sec. III.2, where the emitted secondary electrons are modeled with the Vaughan energy spectrum. Identifying these bands provides insight into multipacting onset and informs suppression strategies.
@@ III Dark Current Simulations
To investigate dark current behavior under short-pulse RF excitation, we perform numerical simulations using the cst Particle-In-Cell (PIC) and particle tracking solvers. This complements the steady-state analytical model by capturing transient effects in realistic cavity geometries. We begin by modeling field emission from high-gradient regions, such as the cavity iris and the photocathode surface. These field-emitted electrons can act as seed particles for multipacting-driven growth and subsequent dark current buildup, as discussed later.
Figure 8: Simulated electron trajectories emitted from (a) the iris and (b) the photocathode surface, under a peak axial electric field amplitude of 325 MV/m at the center of the photocathode surface. Color-coded lines represent trajectories of electrons launched from different initial positions. Emitted current (red, left axis) and downstream current at the end of the beam pipe (blue, right axis; note the different scales) are shown in (c) for iris emission and in (d) for photocathode emission, assuming an estimated average field enhancement factor $\beta_{\mathrm{avg}}=7$.
Figure 9: Temperature rise (right axis) due to field-emission-induced Joule heating, together with the local RF electric field amplitude (left axis). Results are shown for the 9 ns input RF pulse of Fig. 1(c) and for a long pulse with the peak field extended by an additional 100 ns. Calculations assume a field emitter with a radius of 100 nm in a uniform RF field.
@@ III.1 Field Emission Simulations
Field emission is a critical process in RF cavities, where strong local electric fields extract electrons from metal surfaces. These electrons may disrupt the primary beam [50, 51], trigger RF breakdown [29], or initiate dark current through multipacting and other processes. Extensive efforts have been made to model and characterize field emission across various RF structures [31, 52, 27, 53].
Field emission simulations were performed using the cst PIC solver to interpret the low dark current observed during RF conditioning of the $X$-band photogun cavities. As discussed in Sec. I, no dark current signal was detected above the noise threshold [5]. This can be attributed to two factors: (1) the short RF pulse limits the total emitted charge, and (2) emitted electrons must traverse a long, narrow beam pipe (the inner conductor of the coaxial coupler) to reach diagnostics [54]. We illustrate the trajectories of field-emitted electrons in Fig. 8: in (a), most electrons emitted from the iris strike nearby surfaces; in (b), only electrons emitted near the center of the photocathode and within a narrow angular range are able to exit the beam pipe. To quantify dark current transmission, Figs. 8(c) and (d) plot the field-emitted current at the source and the downstream current measured at the end of the beam pipe for emission at the iris and at the photocathode, respectively. In these simulations the RF power is turned on at $t=0$. As noted earlier, no transmitted dark current was detected experimentally [5], so we use this non-detection (implying a per-pulse transmitted charge $1~\text{pC}$ from all dark-current mechanisms) as an upper bound on $\beta$. Field emission simulations indicate that an average $\beta_{\mathrm{avg}}\approx 8$ would yield a transmitted dark-current charge of $1~\text{pC}$ (field emission only); therefore, we adopt $\beta_{\mathrm{avg}}=7$ as the representative case for the simulations in this section. This $\beta_{\mathrm{avg}}$ choice when the entire surface is assumed to emit is equivalent, for the same peak field-emission current, to $\beta\approx 33.5$ assuming an active emitting area of a fraction of $10^{-9}$ of the surface.
Field emission can induce Joule heating, but the shorter emission window under short-pulse operation reduces energy deposition. We model a circular copper emitter with a radius of 100 nm and a uniform local RF electric field with a peak amplitude $\beta E=10$ GV/m. The heating power density is calculated as $j_{F}^{2}/\sigma$, where $j_{F}$ is the Fowler–Nordheim field-emission current density and $\sigma$ is the electrical conductivity. We estimate the temperature rise using a semi-infinite copper half space [55]:
$\color[rgb]{0,0,0}\Delta T(0,t)=\frac{\sqrt{\alpha}}{\kappa\sqrt{\pi}}\int_{0}^{t}\frac{Q(\tau)}{\sqrt{t-\tau}}\,\mathrm{d}\tau,$} (9)
where $\kappa$, $\rho$, and $c_{p}$ are the thermal conductivity, mass density, and specific heat, respectively; $\alpha=\kappa/(\rho c_{p})$ is the thermal diffusivity; and $Q(\tau)$ is the heat flux at time $\tau$. Figure 9 compares the temperature rise in the short- and long-pulse cases. At the same peak field, the short pulse produces a distinctly lower $\Delta T$ than the long flat-top pulse. Therefore, short-pulse operation, particularly with a short flat top, mitigates field-emission-induced heating and increases the thermal margin for high-gradient operation.
@@
Field emission can also contribute to beam loading, but for experimentally relevant $\beta_{\mathrm{avg}}$ values, its direct effect is negligible, as verified with cst PIC simulations. It may still seed multipacting, which can then lead to loading. Field emission transmission simulations in Fig. 8 show that only about $10^{-5}$ of the total electrons emitted from the iris reach the end of the beam pipe. The electrons that strike cavity surfaces may initiate multipacting when resonance conditions are met.
Figure 11: Sample multipacting resonance trajectories simulated in the full cell (defined in Fig. 10). The corresponding axial field amplitudes $E_{z0}$ for resonance orders $N=1\text{--}7$, evaluated at point $P_{3}$ (the center of the full cell), are 106, 61, 39, 31, 26, 22, and 19 MV/m, respectively. These trajectories show good agreement with the analytical results in Fig. 6, as the RF field distribution in the full cell is well approximated by the analytical model.
Figure 12: Sample multipacting resonance trajectories in the half-cell for (a) Region I and (b) Region II. In Region I, the axial fields $E_{z0}$ evaluated at point $P_{1}$ for resonance orders $N=1\text{--}4$ are 112, 61, 39, and 28 MV/m, respectively. In Region II, the axial fields $E_{z0}$ evaluated at point $P_{2}$ for resonance orders $N=1\text{--}5$ are 95, 46, 30, 25, and 21 MV/m, respectively.
III.2 Multipacting simulations
@@
In particle tracking simulations, electron trajectories were analyzed for emission from the sidewalls of the three regions as defined in Fig. 10, the full cell, and Region I and Region II in the half cell. Sample trajectories satisfying multipacting resonance conditions are presented in Figs. 11 and 12. The resonant trajectories and the associated electric field amplitudes in the full cell (Fig. 11) exhibit good agreement with the analytical results shown in Fig. 6 due to the similarity in the RF field pattern, providing further validation of the analytical model.
Figure 13: Temporal evolution of the electric field amplitudes and the number of multipacting electrons for simulations with the secondary electron emission surface defined in (a) Region I, (b) Region II, and (c) the full cell. In each subplot, the on-axis electric field at point $P_{1}$, $P_{2}$, or $P_{3}$ (red solid line), the electric field across the photocathode surface (black dashed line), and the number of multipacting electrons (blue solid line) are shown. For Region I, the dark current peaks corresponding to $N=1\text{--}4$ appear at 113, 66, 43, and 30 MV/m; for Region II, the peaks for $N=1,2,4,5$ are at 100, 47, 28, and 20 MV/m; and for the full cell, the peaks for $N=1,2,3,5$ are at 113, 51, 35, and 27 MV/m. Higher-order resonance modes ($N\geq 6$) also contribute to later peaks in the full cell.
@@
Figure 15: Temporal evolution of the electric field (red) and number of multipacting electrons (blue) using a longer RF pulse with a 30 ns rise from 0 to 200 MW. The simulation was terminated when the multipacting electron count exhausted available memory.
In PIC simulations, we use again the short input RF pulse with a 3 ns rise, 3 ns flat top, and 3 ns fall and a peak power of 200 MW. Seed electrons were introduced at a rate of 400 electrons per RF cycle for 250 consecutive RF cycles, using a procedure similar to those described in described in Refs. [38, 56]. When these seed electrons collide with a surface without multipacting properties, they are removed from the simulation domain upon impact. When they collide with a surface designated for multipacting, secondary emission is modeled using the Vaughan formalism [49], with a maximum yield $\delta_{\text{max}}=2.1$ at an impact energy $W_{i}=150$ eV. The emitted-energy spectrum is modeled with the cst copper parameterization [42]. Figure 13 shows the temporal evolution of the electric field and the number of multipacting electrons in the three regions, where the secondary emission surface is on the sidewalls of Region I, Region II, or the full cell, respectively. The reported count is the total number of multipacting electrons in the simulation domain. Peaks in the electron population correspond to distinct multipacting resonance orders, and the gradient at resonance are in strong agreement with the multipactor resonance trajectory analysis, confirming that electron growth is driven by multipacting resonances at the cavity sidewalls. Under the same input RF pulse, multipacting simulations at the iris and the photocathode surface show no sustained resonant growth of dark electrons.
To further confirm that dark current growth arises from multipacting resonances, we varied the peak power of the input RF pulses while maintaining a fixed temporal profile. The resulting time evolution of the number of multipacting electrons is shown in Fig. 14. For a given cavity geometry, the field gradients at which multipacting resonance occurs are fixed. Adjusting the input power changes the time at which these gradients are reached, thereby shifting the onset of the resonance modes. The dark current peaks consistently occur at the same axial field amplitudes across different power levels, matching the resonance orders identified earlier. Each mode persists for a duration that depends on the field ramp rate. At lower gradients, the field ramp rate is reduced, so the peak dark current is higher due to the longer time window.
Short RF pulses can reduce multipacting dark current, with short rise and fall times being particularly effective in limiting the time window for multipacting resonances. Figure 15 shows the simulation results with longer RF pulses as a comparison, where the gradual rise in field amplitude allows resonance conditions to persist longer, resulting in rapid growth of the multipacting electron population. In contrast, short RF pulses with short rise and fall times sweep through each resonance condition more quickly, reducing the duration over which the field remains within the amplitude window required for sustained multipacting and thus suppressing the buildup of secondary electrons. Therefore, the short-pulse regime reduces dark current by limiting the time available for resonance-driven electron multipacting.
@@ IV Impact of dark electrons
As dark electrons accumulate through field emission and multipacting, electron clouds can form. To observe this process, we perform cst PIC simulations with the cavity operated at the multipacting resonance conditions identified above. Seed electrons are briefly introduced near the cavity sidewall to initiate secondary emission. When the field satisfies resonance conditions, the electron population grows exponentially, forming a localized cloud near the sidewall.
Figure 16: Reflected RF power as a function of time for different plasma electron densities $n_{e}$ and the corresponding plasma frequencies $\omega_{p}$, with the plasma modeled as a 0.2 mm thick layer on the cavity sidewall. All cases use the same short RF input pulse, and the traces are normalized to the no-plasma case.
A critical consequence of multipacting-induced electron clouds is dark-current-induced beam loading. To model this effect, we approximate the electron cloud as a cold plasma layer with a thickness of 0.2 mm, located along the cavity sidewall. The thickness is set to the average maximum wall-normal excursion of the multipacting trajectories (see Figs. 11 and 12). Given the short pulse duration considered here, ion motion is negligible, with the ion velocity estimated to be on the order of $10^{4}$ m/s [57, 58]. The plasma response in this thin layer is described using the Drude dispersion, appropriate for an electron-only, collisionless medium. The effective permittivity of the layer is determined by an assumed plasma electron density $n_{e}$. Figure 16 shows the normalized reflected RF power for different plasma densities, where the corresponding plasma frequency $\omega_{p}=\sqrt{n_{e}e^{2}/(\varepsilon_{0}m)}$ is also displayed ($m$ is the electron mass). As $n_{e}$ increases, the reflected power decreases due to greater energy absorption by the plasma [59]. This behavior is consistent with the experimentally observed drop in reflected power during conditioning of the $X$-band photocathode cavities (see Fig. 2).
@@ V Conclusion and discussion
Recent advances in short-pulse acceleration have opened new opportunities for achieving ultrahigh gradients, but they also require examination of the underlying RF breakdown mechanisms on previously underexplored timescales. Traditional breakdown studies, grounded in long-pulse operation and steady-state field assumptions, fall short in capturing the transient physics characteristic of nanosecond-long RF pulses.
In this work, we use analytical and numerical methods to show that RF pulses on the order of a few nanoseconds can effectively suppress dark current growth by reducing contributions from mechanisms such as field emission, electron multipacting, and plasma formation. A short flat-top duration is essential for mitigating the impact of field emission by reducing the total emitted charge and associated effects such as field-emission-induced Joule heating. For multipacting, our analytical and numerical models identify resonance bands, while time-domain simulations reveal how rapid field ramping shortens the windows for sustained development of multipacting resonances. Short ramp-up and ramp-down times are therefore crucial for suppressing multipacting by limiting the time available for resonance modes to develop. The reduced electron population also lowers the plasma density, mitigating its impact on the RF pulse. These results provide strong evidence that short-pulse operation alters the onset dynamics of dark current and RF breakdown. Our simulation results are also consistent with experimental observations of low dark current and beam loading in the $X$-band photogun cavities at AWA.
To advance the understanding of RF breakdown in the short-pulse regime, future work should focus on modeling the dynamics of plasma formation and its interaction with time-varying RF fields, ideally using first-principles simulation tools. Conventional breakdown models often treat plasma evolution as a multi-stage process, with nanosecond-scale electron dynamics coupled to slower changes in surface morphology. However, in the short-pulse regime, where the electromagnetic fields vary rapidly in time, these transient interactions may dominate the breakdown initiation process. In particular, a self-consistent model that includes both field emission and multipacting, with high-performance computing resources, would enable direct comparison with the experimental setup and is a valuable direction. Investigating these fast-timescale effects will be essential for refining our understanding of the coupled physics in RF breakdown, and short-pulse experiments provide a promising platform for isolating and studying these mechanisms.